# EvolMPNN: Predicting Mutational Effect on Homologous Proteins by Evolution Encoding

### v1

@@ Abstract
Predicting protein properties is paramount for biological and medical advancements. Current protein engineering mutates on a typical protein, called the wild-type, to construct a family of homologous proteins and study their properties. Yet, existing methods easily neglect subtle mutations, failing to capture the effect on the protein properties. To this end, we propose EvolMPNN, Evolution-aware Message Passing Neural Network, to learn evolution-aware protein embeddings. EvolMPNN samples sets of anchor proteins, computes evolutionary information by means of residues and employs a differentiable evolution-aware aggregation scheme over these sampled anchors. This way EvolMPNN can capture the mutation effect on proteins with respect to the anchor proteins. Afterwards, the aggregated evolution-aware embeddings are integrated with sequence embeddings to generate final comprehensive protein embeddings. Our model shows up to $6.4\%$ better than state-of-the-art methods and attains 36×36$\times$36 × inference speedup in comparison with large pre-trained models.
@@ 1 Introduction
Can we predict important properties of a protein by directly observing only the effect of a few mutations on such properties? This basic biological question [1, 2] has recently engaged the machine learning community due to the current availability of benchmark data [3, 4, 5]. Proteins are sequences of amino-acids (residues), which are the cornerstone of life and influence a number of metabolic processes, including diseases [6, 7]. For this reason, protein engineering stands at the forefront of modern biotechnology, offering a remarkable toolkit to manipulate and optimise existing proteins for a wide range of applications, from drug development to personalised therapy [8, 9, 10].
One fundamental process in protein engineering progressively mutates an initial protein, called the wild-type, to study the effect on the protein’s properties [11]. These mutations form a family of homologous proteins as in Figure 1. This process is appealing due to its cheaper cost compared to other methods and reduced time and risk [12, 13].
Yet, the way mutations affect the protein’s properties is not completely understood [14, 15, 16], as it depends on a number of chemical reactions and bonds among residues. For this reason, machine learning offers a viable alternative to model complex interactions among residues. Initial approaches employed feature engineering to capture protein’s evolution [17, 18]; yet, a manual approach is expensive and does not offer enough versatility. Advances in NLP and CV inspired the design of deep protein sequence encoders [19, 20, 21] and general purpose Protein Language Models (PLMs) that are pre-trained on large scale datasets of sequences. Notable PLMs include ProtBert [22], AlphaFold [23], TAPE Transformer [3] and ESM [24]. These models mainly rely on Multiple Sequence Alignments (MSAs) [25] to search on large databases of protein evolution. Nevertheless, this search process is insensitive to subtle yet crucial mutations and introduces additional computational burdens [26, 27].
To overcome the limitations of previous models, we propose EvolMPNN, Evolution-aware Message Passing Neural Network, to predict the mutational effect on homologous proteins. Our fundamental assumption is that there are inherent correlations between protein properties and the sequence differences among them, as shown in Figure 1-(b). EvolMPNN integrates both protein sequence and evolutionary information by identifying where and which mutations occur on the target protein sequence, compared with known protein sequences and predicts the mutational effect on the target protein property. To avoid the costly quadratic pairwise comparison among proteins, we devise a theoretically grounded (see Section 4.6) linear sampling strategy to compute differences only among the proteins and a fixed number of anchor proteins (Section 4.2). We additionally introduce two extensions of our model, EvolGNN and EvolFormer, to include available data on the relation among proteins (Section 4.5). The theoretical computation complexity of proposed methods are provided to guarantee their efficiency and practicality. We apply the proposed methods to three benchmark homologous protein property prediction datasets with nine splits. Empirical evaluation results (Section 5.1) show up to $6.7\%$ Spearman’s $\rho$ correlation improvement over the best performing baseline models, reducing the inference time by 36×36$\times$36 × compared with pre-trained PLMs.
@@
Figure 1: Protein property prediction on homologous protein family. (a) An example homologous protein family with labelled nearby mutants with few mutations. The task is to predict the label of unknown mutants with more mutations. (b) The evolutionary pattern for (a); $(p_{1},m_{1})$ indicates the mutation $m_{1}$ at the $p_{1}$-th position of the protein sequence.
@@ 2 Preliminary and Problem
In the protein engineering process, we first receive a set of proteins =$\mathcal{M}=\{\mathcal{P}_{i}\}_{i=1,2,\dots,M}$ in which each protein can be associated with a label vector $\mathbf{Y}_{i}\in\mathbb{R}^{\theta}$ that describes its biomedical properties, e.g., fitness, stability, fluorescence, solubility, etc. Each protein $\mathcal{P}_{i}$ is a linear chain of amino-acids $\mathcal{P}_{i}=\{r_{j}\}_{j=1,2,\dots,N}$. While a protein sequence folds into specific 3D conformation to perform some biomedical functions, each amino-acid is considered as a residue. Residues are connected to one another by peptide bonds and can interact with each other by different chemical bounds [6]. In short, the function of a protein is mainly determined by the chemical interactions between residues. Since the 3D structure is missing in benchmark datasets [3, 4, 5], we assume no 3D protein information in this paper.
@@
Research Problem. Protein engineering based on homologous proteins is a promising and essential direction for designing novel proteins of desired properties [29, 30]. Understanding the relation between protein sequence and property is one essential step. Practically, biologists perform experiments in the lab to label the property $\mathbf{Y}_{\textsc{Train}}$ of a set of protein $\mathcal{M}_{\textsc{Train}}\subset\mathcal{M}$ and the follow-up task is predicting $\hat{\mathbf{Y}}_{\textsc{Test}}$ of the rest proteins $\mathcal{M}_{\textsc{Test}}\subset\mathcal{M}$. However, homologous proteins typically have similarities in their amino-acid sequences, structures, and functions due to their shared ancestry. Accurately predicting the homologous protein property by distinguishing these subtle yet crucial differences is still an open challenge.
@@ 3 Related Work
Feature Engineering. Besides conducting manual experiments in labs to measure protein properties, the basic solution is to design different feature engineering methods based on relevant biological knowledge, to extract useful information from protein sequence [31, 18, 32]. [4] introduce using Levenshtein distance [33] and BLOSUM62-score [34] relative to wild-type to design protein sequence features. In another benchmark work, [5] adopt another two typical protein sequence feature descriptors, i.e., Dipeptide Deviation from Expected Mean (DDE) [17] and Moran correlation (Moran) [18]. For more engineering methods, refer to the comprehensive review [35].
Protein Representation Learning. In the last decades, empowered by the outstanding achievements of machine learning and deep learning, protein representation learning has revolutionised protein property prediction research. Early work along this line adopts the idea of word2vec [36] to protein sequences [37, 38]. To increase model capacity, deeper protein sequence encoders were proposed by the Computer Vision (CV) and Nature Language Processing (NLP) communities [19, 20, 21]. The latest works develop Protein Language Models (PLMs), which focus on employing deep sequence encoder models for protein sequences and are pre-trained on million- or billion-scale sequences. Well-known works include ProtBert [22], AlphaFold [23], TAPE Transformer [3] and ESM [24]. However, most existing work does not pay enough attention to these subtle yet crucial differences in homologous proteins. [24, 23] explore protein Multiple Sequence Alignments (MSAs) [39, 25] to capture the mutational effect. Nevertheless, the MSA searching process introduces additional computational burden and is insensitive to subtle but crucial sequence differences [26]. [27] indicate the shortcomings of MSAs on easily neglecting the presence of minor mutations, which can propagate errors to downstream protein sequence representation learning tasks. This paper also lies in this direction, we propose a novel solution for the challenging homologous protein property prediction tasks.
@@ 4.1 Embedding Initialisation
Protein Sequence Embedding. Given a set of proteins =$\mathcal{M}=\{\mathcal{P}_{i}\}_{i=1,2,\dots,M}$, we first adopt a (parameter-frozen) PLM model [40, 25] [^1] We do not fine-tune PLM in this paper for efficiency consideration. as protein sequence encoder to initialise the protein sequence embedding ($\mathbf{H}$) for every protein $\mathcal{P}_{i}$, which include macro (i.e., protein sequence) level information as the primary embedding.
@@
Residue Embedding. In order to properly capture the evolution information in homologous proteins, we delve into the residue level for micro clues. We adopt two residue embedding initialisation approaches, i.e., one-hot encoding ($\Phi^{\textsc{OH}}$) and pre-trained PLM encoder ($\Phi^{\textsc{PLM}}$), to generate protein’s initial residue embeddings $\mathbf{X}_{i}=\{\mathbf{x}^{i}_{j}\}_{j=1,2,\dots,N}$, where $\mathbf{x}^{i}_{j}\in\mathbb{R}^{d}$. In particular, $\Phi^{\textsc{OH}}$ assigns each protein residue [^2] There are $20$ different amino-acid residues commonly found in proteins with a binary feature vector $\mathbf{x}^{i}_{j}$, where $\mathbf{x}^{i}_{jb}=1$ indicates the appearance of the $b$-th residue at $\mathcal{P}_{i}$’s $j$-th position. By stacking $N$ residues’ feature vectors into a matrix, we can obtain $\mathbf{X}_{i}\in\mathbb{R}^{N\times d}$. On the other hand, following the benchmark implementations [41], PlmEncoder can export residue embeddings similar to Eq. 1. Formally, $\Phi^{\textsc{PLM}}$ initialises protein residue embeddings as $\mathbf{X}_{i}=\textsc{PlmEncoder}(\{r_{j}\}_{j=1,2,\dots,N})$.
Position Embedding. Another essential component of existing PLM is the positional encoding, which was first proposed by [21]. This positional encoding effectively captures the relative structural information between entities and integrates it with the model [42]. In our case, correctly recording the position of each residue in the protein sequence plays an essential role in identifying each protein’s corresponding mutations. Because the same mutation that occurs at different positions may lead to disparate influences on protein property. Therefore, after initialising residue embeddings, we further apply positional embedding on each protein’s residues. We adopt a methodology that reminisces [43] that demonstrates the paramount importance of assigning each residue with a unique position embedding. As such, we randomly initialise a set of $d$ position embeddings $\Phi^{\textsc{Pos}}\in\mathbb{R}^{N\times d}$. We denote the residue embedding empowered by position embedding as $\hat{\mathbf{X}}_{i}=\mathbf{X}_{i}\odot\Phi^{\textsc{Pos}}$.
@@ 4.2 Residue Embedding Update
The 3D protein folding depends on the strength of different chemical bonds between residues to maintain a stable 3D structure. Previous studies carefully designed residue contact maps to model the residue-residue interactions to learn effective residue embeddings [40, 44]. In this paper, we adopt the residue-residue interaction to update residue embeddings but eschew the requirement of manually designing the contact map. Instead, we assume the existence of an implicit fully connected residue contact map of each protein $\mathcal{P}_{i}$ and implement the Transformer model [21, 45] to adaptively update residue embeddings. Denote $\mathbf{R}^{(\ell)}_{i}$ as the input to the $(\ell+1)$-th layer, with the first $\mathbf{R}^{(0)}_{i}=\hat{\mathbf{X}}_{i}$ be the input encoding. The $(\ell+1)$-th layer of residue embedding update module can be formally defined as follows:
@@
where $\mathbf{W}^{\ell,h}_{O}\in\mathbb{R}^{d_{H}\times d}$, $\mathbf{W}^{l,h}_{Q}$, $\mathbf{W}^{l,h}_{K}$, $\mathbf{W}^{l,h}_{V}\in\mathbb{R}^{d\times d_{H}}$, $\mathbf{W}^{\ell}_{1}\in\mathbb{R}^{d\times r}$, $\mathbf{W}^{\ell}_{2}\in\mathbb{R}^{d_{t}\times d}$, $H$ is the number of attention heads, $d_{H}$ is the dimension of each head, $d_{t}$ is the dimension of the hidden layer, ELU [46] is an activation function, and $\mathbf{Att}^{h}_{i}(\mathbf{R}^{(\ell)}_{i})$ refers to as the attention matrix. After each Transformer layer, we add a normalisation layer i.e., LayerNorm [47], to reduce the over-fitting problem proposed by [21]. After stacking $L_{r}$ layers, we obtain the final residue embeddings as $\mathbf{R}_{i}=\mathbf{R}^{(L_{r})}_{i}$.
@@ 4.3 Evolution Encoding
Assumption 1 (Protein Property Relevance).
@@
To address the aforementioned problems, we propose an anchor-based protein evolution encoding method. Specifically, denote $\mathbf{H}^{(\ell)}_{i}$ as the input to the $(\ell+1)$-th block and define $\mathbf{H}^{(0)}_{i}=\mathbf{H}_{i}$. The evolution localisation encoding of the $(\ell+1)$-th layer contains the following key components: (i) $k$ anchor protein $\{\mathcal{P}_{S_{i}}\}_{i=1,2,\dots,k}$ selection. (ii) Evolutionary information encoding function $\mathrm{F}_{\textsc{Diff}}$ that computes the difference between residues of each protein and those of the anchor protein, and target protein’s evolutionary information is generated by summarising the obtained differences:
@@
where Combine can be implemented as differentiable operators, such as, Concatenate, MAX Pool MEAN Pool and SUM Pool; here we use the MEAN Pool to obtain $\mathbf{d}_{ij}\in\mathbb{R}^{d}$. (iii) Message computation function $\mathrm{F}_{\textsc{Message}}$ that combines protein sequence feature information of two proteins with their evolutionary differences. We empirically find that the simple element-wise product between $\mathbf{H}^{(\ell)}_{j}$ and $\mathbf{d}_{ij}$ attains good results
@@
(iv) Aggregate messages from $k$ anchors and combine them with protein’s embedding as follow:

### v2

# Efficiently Predicting Mutational Effect on Homologous Proteins by Evolution Encoding

@@ Abstract
Predicting protein properties is paramount for biological and medical advancements. Current protein engineering mutates on a typical protein, called the wild-type, to construct a family of homologous proteins and study their properties. Yet, existing methods easily neglect subtle mutations, failing to capture the effect on the protein properties. To this end, we propose EvolMPNN, Evolution-aware Message Passing Neural Network, an efficient model to learn evolution-aware protein embeddings. EvolMPNN samples sets of anchor proteins, computes evolutionary information by means of residues and employs a differentiable evolution-aware aggregation scheme over these sampled anchors. This way, EvolMPNN can efficiently utilise a novel message-passing method to capture the mutation effect on proteins with respect to the anchor proteins. Afterwards, the aggregated evolution-aware embeddings are integrated with sequence embeddings to generate final comprehensive protein embeddings. Our model shows up to $6.4\%$ better than state-of-the-art methods and attains 36×36$\times$36 × inference speedup in comparison with large pre-trained models. Code and models are available at https://github.com/zhiqiangzhongddu/EvolMPNN.
@@ 1 Introduction
Can we predict important properties of a protein by directly observing only the effect of a few mutations on such properties? This basic biological question [16] has recently engaged the machine learning community due to the current availability of benchmark data [30, 8, 44]. Proteins are sequences of amino acids (or residues), which are the cornerstone of life and influence a number of metabolic processes, including diseases [28]. For this reason, protein engineering stands at the forefront of modern biotechnology, offering a remarkable toolkit to manipulate and optimise existing proteins for a wide range of applications, from drug development to personalised therapy [1].
@@
Figure 1: Protein property prediction on homologous protein family. (a) An example homologous protein family with labelled nearby mutants with few mutations. We aim to predict the label of unknown mutants with more mutations. (b) The evolutionary pattern for (a). For instance, $\mathbf{Y}_{0}$ is the label vector of the corresponding protein sequence, and $(p_{1},m_{1})$ indicates mutation $m_{1}$ at position $p_{1}$ of the protein’s amino acid sequence.
One fundamental process in protein engineering progressively mutates an initial protein, called the wild-type, to study the effect on the protein’s properties [38]. These mutations form a family of homologous proteins as in Figure 1. This process is appealing due to its cheaper cost compared to other methods and reduced time and risk [41].
Yet, the way mutations affect the protein’s properties is not completely understood [5], as it depends on a number of chemical reactions and bonds among residues. For this reason, machine learning offers a viable alternative to model complex interactions among residues. Initial approaches employed feature engineering to capture protein’s evolution [35, 14]; yet, a manual approach is expensive and does not offer enough versatility. Advances in NLP and CV inspired the design of deep protein sequence encoders [18, 39] and general purpose Protein Language Models (PLMs) that are pre-trained on large-scale datasets of sequences. Notable PLMs include ProtBert [4], AlphaFold [19], TAPE Transformer [30] and ESM [33]. These models mainly rely on Multiple Sequence Alignments (MSAs) [25] to search on large databases of protein evolution. While this process focuses on conserved regions, it is insensitive to subtle yet crucial mutations in less conserved regions and introduces additional computational burdens [29, 6].
To overcome the limitations of previous models, we propose EvolMPNN, Evolution-aware Message Passing Neural Network, to predict the mutational effect on homologous proteins. Our fundamental assumption is that there are inherent correlations between protein properties and the sequence differences among them, as shown in Figure 1-(b). EvolMPNN devises a novel message-passing method to integrate both protein sequence and evolutionary information by identifying where and which mutations occur on the target protein sequence, compared with known protein sequences and predicts the mutational effect on the target protein property. To avoid the costly quadratic pairwise comparison among proteins, we devise a theoretically grounded (see Section 4.6) linear sampling strategy to compute differences only among the proteins and a fixed number of anchor proteins (Section 4.2). We additionally introduce two extensions of our model, EvolGNN and EvolFormer, to include available data on the relation among proteins (Section 4.5). The theoretical computation complexities of proposed methods are provided to guarantee their efficiency and practicality. We apply the proposed methods to three benchmark homologous protein property prediction datasets with nine splits. Empirical evaluation results (Section 5.1) show up to $6.7\%$ Spearman’s $\rho$ correlation improvement over the best performing baseline models, reducing the inference time by 36×36$\times$36 × compared with pre-trained PLMs.
@@ 2 Preliminary and Problem
In protein engineering, we first receive a set of proteins =$\mathcal{M}=\{\mathcal{P}_{i}\}_{i=1,2,\dots,M}$ in which each protein can be associated with a label vector $\mathbf{Y}_{i}\in\mathbb{R}^{\theta}$ that describes its biomedical properties, e.g., fitness, stability, fluorescence, and solubility. Each protein $\mathcal{P}_{i}$ is a linear chain of amino-acids $\mathcal{P}_{i}=\{r_{j}\}_{j=1,2,\dots,N}$. While a protein sequence folds into specific 3D conformation to perform some biomedical functions, each amino-acid is considered as a residue. Residues form peptide bonds and can interact with each other by different chemical bounds [28]. In short, the function of a protein is mainly determined by the chemical interactions between residues. Since the 3D structure is missing in benchmark datasets [30, 8, 44], we assume no 3D protein information in this paper.
@@
Research Problem. Protein engineering based on homologous proteins is a promising and essential direction for designing novel proteins of desired properties [16]. Understanding the relation between protein sequence and property is one essential step. Practically, biologists perform experiments in the lab to manually label the property $\mathbf{Y}_{\textsc{Train}}$ of a set of protein $\mathcal{M}_{\textsc{Train}}\subset\mathcal{M}$ and the follow-up task is predicting $\hat{\mathbf{Y}}_{\textsc{Test}}$ of the rest proteins $\mathcal{M}_{\textsc{Test}}\subset\mathcal{M}$. However, due to their shared ancestry, homologous proteins typically have similarities in their amino acid sequences, structures, and functions. Accurately predicting the homologous protein property by distinguishing these subtle yet crucial differences is still an open challenge.
@@ 3 Related Work
Feature Engineering. Besides conducting manual experiments in labs to measure protein properties, the basic solution is to design different feature engineering methods based on relevant biological knowledge, to extract useful information from protein sequence [14]. [8] introduce using Levenshtein distance [22] and BLOSUM62-score [10] relative to wild-type to design protein sequence features. In another benchmark work, [44] adopt another two typical protein sequence feature descriptors, i.e., Dipeptide Deviation from Expected Mean (DDE) [35] and Moran correlation (Moran) [14]. For more engineering methods, refer to the comprehensive review [21].
Protein Representation Learning. In the last decades, propelled by the outstanding achievements of machine learning and deep learning, protein representation learning has revolutionised protein property prediction research. Early work along this line adopts the idea of word2vec [26] to protein sequences. To increase model capacity, deeper protein sequence encoders were proposed by the Computer Vision and Nature Language Processing communities [18, 39]. The latest works develop Protein Language Models, which focus on employing deep sequence encoder models for protein sequences and are pre-trained on million- or billion-scale sequences. Well-known works include ProtBert [4], AlphaFold [19], TAPE Transformer [30], ESM [33], AlphaFold2 [19] and Ankh [11]. However, most current research on protein sequences defines it as a language problem and does not sufficiently consider the subtle yet crucial evolutionary patterns in homologous proteins. For instance, [33, 19] explore protein Multiple Sequence Alignments (MSAs) [31, 25] to capture the mutational effect. Nevertheless, the MSA searching process introduces additional computational burden and is insensitive to subtle but crucial sequence differences [29]. [6] indicate the shortcomings of MSAs on easily neglecting the presence of minor mutations, which can propagate errors to downstream protein sequence representation learning tasks. The message-passing idea has been explored in protein folding tasks [19, 9]. We build on this direction, by devising a novel model that captures subtleties in the mutations for homologous protein property prediction tasks.
@@ 4.1 Embedding Initialisation
Protein Sequence Embedding. Given a set of proteins =$\mathcal{M}=\{\mathcal{P}_{i}\}_{i=1,2,\dots,M}$, we first adopt a (parameter-frozen) PLM model [32, 25] [^1] We do not fine-tune PLM in this paper for efficiency consideration. as protein sequence encoder to initialise protein-sequence embeddings ($\mathbf{H}$) as one $d$-dimensional real-valued vector for every protein $\mathcal{P}_{i}$, which include macro (i.e., protein sequence) level information as the primary embedding.
@@
Residue Embedding Initialisation. In order to properly capture the evolution information in homologous proteins, we delve into the residue level for micro clues. We adopt two residue embedding initialisation approaches, i.e., one-hot encoding ($\Phi^{\textsc{OH}}$) and pre-trained PLM encoder ($\Phi^{\textsc{PLM}}$), to generate protein’s initial residue embeddings $\mathbf{X}_{i}=\{\mathbf{x}^{i}_{j}\}_{j=1,2,\dots,N}$, where $\mathbf{x}^{i}_{j}\in\mathbb{R}^{d}$. In particular, $\Phi^{\textsc{OH}}$ assigns each protein residue [^2] There are $20$ different amino acid residues commonly found in proteins with a binary feature vector $\mathbf{x}^{i}_{j}$, where $\mathbf{x}^{i}_{jb}=1$ indicates the appearance of the $b$-th residue at $\mathcal{P}_{i}$’s $j$-th position. By stacking $N$ residues’ feature vectors into a matrix, we can obtain $\mathbf{X}_{i}\in\mathbb{R}^{N\times d}$. On the other hand, following the benchmark implementations [48], PlmEncoder can export residue embeddings similar to Eq. 1. Formally, $\Phi^{\textsc{PLM}}$ initialises protein residue embeddings as $\mathbf{X}_{i}=\textsc{PlmEncoder}(\{r_{j}\}_{j=1,2,\dots,N})$.
Position Embedding. Another essential component of existing PLM is the positional encoding, which was first proposed by [39]. This positional encoding effectively captures the relative structural information between residues and integrates it with the model. In our case, correctly recording the position of each residue in the protein sequence plays an essential role in identifying each protein’s corresponding mutations. Because a mutation occurring at different positions can have varying effects on protein properties due to its impact on protein structure. Therefore, after initialising residue embeddings, we further apply positional embedding on each protein’s residues. Particularly, we randomly initialise a set of $d$ position embeddings $\Phi^{\textsc{Pos}}\in\mathbb{R}^{N\times d}$, and it will be learned in the training process of the entire framework. We denote the residue embedding empowered by position embedding as $\hat{\mathbf{X}}_{i}=\mathbf{X}_{i}\odot\Phi^{\textsc{Pos}}$, where $\odot$ indicates the element-wise multiplication.
@@ 4.2 Residue Embedding Update
To maintain a stable 3D structure, 3D protein folding depends on the strength of different chemical bonds between residues. Previous studies manually designed residue contact maps to model the residue-residue interactions to learn effective residue embeddings [32, 17]. In this paper, we adopt the residue-residue interaction to update residue embeddings but eschew the requirement of manually designing the contact map. Instead, we assume the existence of an implicit fully connected residue contact map of each protein $\mathcal{P}_{i}$ and implement the Transformer model [39, 43] to adaptively update residue embeddings. Denote $\mathbf{R}^{(\ell)}_{i}$ as the input to the $(\ell+1)$-th layer, with the first $\mathbf{R}^{(0)}_{i}=\hat{\mathbf{X}}_{i}$ be the input encoding. The $(\ell+1)$-th layer of residue embedding update module can be formally defined as follows:
@@
where $\mathbf{W}^{\ell,h}_{O}\in\mathbb{R}^{d_{H}\times d}$, $\mathbf{W}^{l,h}_{Q}$, $\mathbf{W}^{l,h}_{K}$, $\mathbf{W}^{l,h}_{V}\in\mathbb{R}^{d\times d_{H}}$, $\mathbf{W}^{\ell}_{1}\in\mathbb{R}^{d\times r}$, $\mathbf{W}^{\ell}_{2}\in\mathbb{R}^{d_{t}\times d}$ are learnable parameters, $H$ is the number of attention heads, $d_{H}$ is the dimension of each head, $d_{t}$ is the dimension of the hidden layer, ELU [7] is an activation function, and $\mathbf{Att}^{h}_{i}(\mathbf{R}^{(\ell)}_{i})$ refers to as the attention matrix. After each Transformer layer, we add a normalisation layer i.e., LayerNorm [2], to reduce the over-fitting problem proposed by [39]. After stacking $L_{r}$ layers, we obtain the final residue embeddings as $\mathbf{R}_{i}=\mathbf{R}^{(L_{r})}_{i}$.
@@ 4.3 Evolution Encoding
Assumption 1 (Protein Property Relevance)
@@
To address the aforementioned problems, instead of constructing the phylogenetic tree, we compute the distance among a few sampled proteins we call anchor proteins and all the other proteins. Theoretical analysis to validate this design is discussed in Section 4.6. Specifically, denote $\mathbf{H}^{(\ell)}_{i}$ as the input to the $(\ell+1)$-th block and define $\mathbf{H}^{(0)}_{i}=\mathbf{H}_{i}$. The evolution localisation encoding of the $(\ell+1)$-th layer contains the following key components: (i) $k$ anchor protein $\{\mathcal{P}_{S_{i}}\}_{i=1,2,\dots,k}$ selection. (ii) Evolutionary information encoding function $\mathrm{F}_{\textsc{Diff}}$ that computes the difference between residues of each protein and those of the anchor protein, and target protein’s evolutionary information is generated by summarising the obtained differences $\mathbf{d}_{ij}$ as follows:
@@
where Combine can be implemented as differentiable operators, such as, Concatenate, MAX Pool MEAN Pool and SUM Pool; here we use the MEAN Pool to obtain $\mathbf{d}_{ij}\in\mathbb{R}^{d}$. (iii) Message computation function $\mathrm{F}_{\textsc{Message}}$ that integrates protein features and evolutionary information as one message from an anchor protein. Specifically, $\mathrm{F}_{\textsc{Message}}$ combines protein sequence feature information of two proteins with their evolutionary differences. We empirically find that the simple element-wise product between $\mathbf{H}^{(\ell)}_{j}$ and $\mathbf{d}_{ij}$ attains good results
@@
(iv) Aggregate evolutionary messages from $k$ anchors and combine them with the protein’s embedding as the updated protein embedding, which contains the protein sequence and evolutionary information: